OpenStudy (anonymous):

Use the rules of exponents to evaluate or simplify. Write without negative exponents. (-243)^(3/5)

OpenStudy (anonymous):

\[-243^{\dfrac{3}{5}}=-1(\sqrt[5]{243})^3=-27\]

OpenStudy (anonymous):

Thanks HomeZone.